# Bounds to the Basset-Boussinesq force on particle laden stratified flows

### v1

@@ IV Numerical simulations
We now focus our attention on the numerical validation of the condition obtained in Sec. III for regimes in which the Basset-Boussinesq force can be neglected or not. To this end we performed several numerical simulations of stably stratified turbulence (see table 1), each of them with different particles (see table 1). The Boussinesq fluid equations given by Eqs. (4) and (5) were numerically solved in a triply periodic domain using a parallelized and fully dealiased pseudo-spectral method, along with a second-order Runge-Kutta scheme for time integration [29]. For the evolution of inertial particles satisfying Eq. (6) we used third-order spline interpolation to estimate forces at the particles positions, and a second-order Runge-Kutta method for time integration [30]. The Basset-Boussinesq force was computed using the second-order method described in van Hinsberg et al. [28].
Numerical simulations were carried out using a spatial resolution of $N_{x}=N_{y}=768$ and $N_{z}=192$ grid points. The domain had dimensions $L_{x}=L_{y}=2\pi L_{0}$ in the horizontal directions, and $L_{z}=H=\pi L_{0}/2$ in the vertical direction, where $L_{0}$ is a unit length. We considered four different Brunt-Väisälä frequencies (see table 1), measured in units of the inverse of a unit turnover time $T_{0}=L_{0}/U_{0}$, with $U_{0}$ representing a unit velocity. For simplicity, all simulations had a Prandtl number $\textrm{Pr}=\nu/\kappa=1$. The kinematic viscosity was chosen such that the Kolmogorov scale $\eta=(\nu^{3}/\varepsilon)^{1/4}\approx 0.005L_{0}$ was well resolved, where the kinetic energy dissipation rate is $\varepsilon=\nu\left<|\bm{\omega}|^{2}\right>$ and $\bm{\omega}=\bm{\nabla}\times{\bf u}$ is the vorticity. This results in $\kappa\eta\approx 1.6$, where $\kappa=N_{x}/(3L_{0})$ is the maximum resolved wave number when using the $2/3$ rule for dealiasing, ensuring spatially well-resolved simulations [31, 32].
@@
Particles were initialized randomly in a horizontal strip of width $H/5$, centered around $z_{0}=H/2$, and at a time in which the flows had reached a turbulent steady state (i.e., after flow integration for over 60 large-scale turnover times). The initial velocities of the particles matched the fluid velocity at the center of each particle. The particles were one-way coupled, essentially functioning as test particles. They neither collide with each other nor contribute to the flow dynamics. To each fluid simulation in Table, four sets of particles were introduced, each containing 25,000 particles, each set characterized by distinct values of $\tau_{p}$. Particles were integrated for more than 15 large-scale turnover times. Moreover, for each set of particles, integrations were done solving Eq. (6) with and without the Basset-Boussinesq history term (labed in the following respectively as “w/H” and “w/oH”). This resulted in a cumulative count of 32 particle datasets, each with their corresponding Reynolds, Froude, and Stokes numbers as defined next.
@@
which quantifies the turbulence intensity at the buoyancy scale $L_{b}=U/N$, and plays a crucial role in describing the flow behavior. When $\textrm{Rb}\gg 1$ turbulence is strong even in the presence of stratification, while for $\textrm{Rb}\ll 1$ turbulent motions are significantly suppressed by viscosity in each stratified layer. Geophysical flows typically exhibit large values of Rb [33], but computational constraints impose severe restrictions on the range of values of Rb that can be simulated. Previous studies indicate that $\textrm{Rb}>10$ is sufficient for the flow to sustain significant turbulence at small scales [34]. Considering computational limitations and the need to explore parameter space, we consider flows with Rb ranging from 6 to 130. Another relevant length scale to characterize the small scales of stratified turbulence is the Ozmidov scale, $L_{Oz}=2\pi/k_{Oz}$, with $k_{Oz}=\sqrt{N^{3}/\varepsilon}$. At scales significantly smaller than $L_{Oz}$ we expect the flow to asymptotically recover isotropy. Thus, when Rb is sufficiently large, and when $L_{Oz}$ is larger than the the Kolmogorov dissipation scale, we can expect small scale turbulence to be stronger, and to affect as a result the particle dynamics and transport.
@@
where St is the Stokes number, $\tau_{\eta}=(\nu/\varepsilon)^{1/2}$ is the Kolmogorov dissipation time, and $\textrm{Re}_{p}$ is the particle Reynolds number.
Tables 1 and 2 provide all these dimensionless numbers and characteristic scales for the simulations. Note these tables should be read together, as we can have, e.g., particles with $\textrm{St}=0.3$ in a flow with $N=4/T_{0}$ (with or without the Basset-Boussinesq history term), or the same particles but in a flow with $N=8/T_{0}$, $12/T_{0}$, or $20/T_{0}$.
@@ V Numerical results
Figure 1: (a) Ratio of the squared infinity norm of the vertical components of the Basset-Boussinesq force to the Stokes force, as a function of the buoyancy Stokes number Sb for all particles and simulations. A linear relation with slope of 1 is shown as a reference. (b) Same for the infinity norms considering all components of both foces.
@@ V.1 Relation between the forces and the buoyancy Stokes number
The first direct test of the bound given by Eq. (18) considers the strength of the forces in all simulations with particles that include the Basset-Boussinesq force. To estimate $\|{\bf F}_{\textrm{BB}}\|_{\infty}$, $\|{\bf F}_{\textrm{BB}}\cdot\hat{z}\|_{\infty}$, $\|{\bf F}_{\textrm{St}}\|_{\infty}$, and $\|{\bf F}_{\textrm{St}}\cdot\hat{z}\|_{\infty}$, the maximum of the vector modulus and of the vertical component of each force were calculated for each particle in the simulations over all times, and then the average values of ${\|{\bf F}_{\textrm{BB}}\|^{2}_{\infty}}/{\|{\bf F}_{\textrm{St}}\|^{2}_{\infty}}$ and of ${\|{\bf F}_{\textrm{BB}}\cdot\hat{z}\|^{2}_{\infty}}/{\|{\bf F}_{\textrm{St}}\cdot\hat{z}\|^{2}_{\infty}}$ were computed for each value of Sb. The results are shown in Fig. 1. Note that ${\|{\bf F}_{\textrm{BB}}\cdot\hat{z}\|^{2}_{\infty}}/{\|{\bf F}_{\textrm{St}}\cdot\hat{z}\|^{2}_{\infty}}<1$ when $\textrm{St}<1$. The condition on ${\|{\bf F}_{\textrm{BB}}\|^{2}_{\infty}}/{\|{\bf F}_{\textrm{Sb}}\|^{2}_{\infty}}$ being small is also bounded by St, confirming that the vertical component of the Basset-Boussinesq force is dominant, and it is the component of interest to obtain a bound on the effect of this force on the particles dynamics.
@@ V.2 Influence of the Basset-Boussinesq force on particle dispersion
where the subindex $i=1,2$ stands for the $x$ or $y$ coordinates, and the supraindex $j$ can be w/H or w/oH (i.e., integration with or without the history term). Figure 2 shows the ratio of the averages over all particles of the mean square horizontal displacements with and without the history term, $\langle{\delta_{x}^{(\textrm{w/H})}}^{2}+{\delta_{y}^{(\textrm{w/H})}}^{2}\rangle/\langle{\delta_{x}^{(\textrm{w/oH})}}^{2}+{\delta_{y}^{(\textrm{w/oH})}}^{2}\rangle$, as a function of time for different Froude and Stokes numbers.
For small times compared with the Brunt-Väisälä period, $2\pi/N$, the ratio remains close to unity, but as time becomes close to $2\pi/N$ differences in the mean squared horizontal dispersion develop. For small values of St (corresponding to small $\tau_{p}$) and large values of Fr (i.e., small values of $N$), the ratio hovers around unity and the differences between the particles w/H and w/oH remain below $4\%$. As the value of St of the particles increases, the ratio deviates from unity. For fixed St values, the ratio also increases as Fr decreases. This increase in the differences w/H and w/oH is compatible with the increase in Sb. For large values of Sb differences in the horizontal dispersion are between $10\%$ to $15\%$, depending on the value of Fr. However, when $\textrm{Sb}<1$, errors when neglecting the Basset-Boussinesq force consistently stay below $5\%$. It is noteworthy that in many cases the largest errors are found for intermediate times. Van Aartrijk and Clercx [25] reported that in stratified flows different regimes develop in the vertical and horizontal dispersion of particles, with the Basset-Boussinesq force having a significant impact in the development and extension of transients at times close to the Brunt-Väisälä period.
@@ V.4 Effect of the Basset-Boussinesq force on particle clustering
The vertical confinement of neutrally buoyant particles in thin layers resulting from the flow stratification has a strong impact on inertial particle clustering [27]. The aggregation of particles is relevant in oceanic flows, in which even the dynamics of large particles are sometimes modeled using modified versions of the Maxey-Riley equation. An example can be found in the study of sargassum, a type of seaweed that serves as the habitat for marine fauna but can also pose environmental challenges due to elevated levels of arsenic and heavy metals when decomposing near coastlines [35, 36]. Another example involves phytoplankton, that aggregates creating intricate structures spanning kilometers [37, 38], and forming thin layers at depths correlated with regions of pronounced fluid density gradients and vertical shear typically occurring near the bottom of the oceanic mixed layer [39]. We therefore quantify in this subsection the effect of considering or neglecting the Basset-Boussinesq force on particle clustering. To quantify the aggregation of particles we use a Voronoï tessellation. This technique has been used before to characterize the preferential concentration of particles in laboratory experiments and in numerical simulations [40, 41, 42, 7, 43, 44, 45, 46].
In a Voronoï tessellation each particle is assigned a cell, corresponding to all the volume (or area, in two-dimensional slices) that is closer to that particle than to any other neighbouring particle. The size of the cells are then inversely proportional to the particle density in that region: larger cells correspond to voids (i.e., regions with far apart particles), while smaller cells correspond to clusters (i.e, regions with particles closer to each other). The standard deviation of the volume (or area) of all the Voronoï cells can then we used to quantify the amount of clustering in the flow [40, 41]: when the standard deviation is larger than that of a random Poisson process, there is an excess of clusters and of voids compared with a random homogeneous distribution of particles. While both three- and two-dimensional tessellations have been used to study particle clustering in homogeneous and isotropic turbulence, in the case of stably stratified flows the confinement of all particles in a layer near a the neutrally buoyant level makes it reasonable to limit the study to 2D tessellations, by projecting all particles into the horizontal plane (see, e.g., [27]).
@@ VI Conclusions
We obtained a bound that is useful to determine under what conditions the Basset-Boussinesq force in the Maxey–Riley equation for inertial particles can be neglected, when the particles are submerged in a stratified flow. This bound motivated the definition of a buoyancy Stokes number, $\textrm{Sb}=N\tau_{p}$ which is the Stokes number of the particles at the fluid buoyancy scale, i.e., it is the ratio of the particle response time to the Brunt-Väisälä period. For sufficiently small Sb the effect of the Basset-Boussinesq force becomes negligible.
This condition was validated using direct numerical simulations of small particles in stably stratified turbulent flows, exploring parameter space by varying the fluid Froude number and the particles Stokes number. Numerical integration of the particles was also performed considering and neglecting the Basset-Boussinesq force in the Maxey–Riley equation, to quantify differences between these two cases.
Using the numerical simulations we computed the infinity norm of the forces acting on the particles, and showed that the infinity norm of the Basset-Boussinesq force becomes smaller than the infinity norm of the Stokes force when $\textrm{Sb}<1$. Three other observables were considered: single particle dispersion (i.e., the mean squared distance traveled by the particles), the mean squared velocity of the particles, and particle clustering. In all cases, differences between simulations with and without the Basset-Boussinesq force were small when $\textrm{Sb}<1$, and grew rapidly with Sb when $\textrm{Sb}>1$. Thus, the Basset-Boussinesq force must be considered in particles with large inertia, or in particle laden fluids with strong stratification. The results presented here allow the estimation of the conditions under which this force becomes relevant in different geophysical contexts [37, 22, 25, 35, 36, 5, 6].

### v2

@@ IV Numerical simulations
We now focus our attention on the numerical validation of the condition obtained in Sec. III for regimes in which the Basset-Boussinesq force can be neglected or not. To this end we performed several numerical simulations of stably stratified turbulence (see table 1), each of them with different particles (see table 2). The Boussinesq fluid equations given by Eqs. (4) and (5) were numerically solved in a triply periodic domain using a parallelized and fully dealiased pseudo-spectral method, along with a second-order Runge-Kutta scheme for time integration [45]. For the evolution of inertial particles satisfying Eq. (6) we used third-order spline interpolation to estimate forces at the particles positions, and a second-order Runge-Kutta method for time integration [46]. The Basset-Boussinesq force was computed using the second-order method described in van Hinsberg et al. [29]. We explicitly verified that errors in this numerical integration scheme for the Basset-Boussinesq force remained small, by comparing against explicit integration of a small fraction of the particles in simulations with the same fluid and particle parameters.
@@
Numerical simulations were carried out using a spatial resolution of $N_{x}=N_{y}=768$ and $N_{z}=192$ grid points. The domain had dimensions $L_{x}=L_{y}=2\pi L_{0}$ in the horizontal directions, and $L_{z}=H=\pi L_{0}/2$ in the vertical direction, where $L_{0}$ is a unit length. We considered four different Brunt-Väisälä frequencies (see table 1), measured in units of the inverse of a unit turnover time $T_{0}=L_{0}/U_{0}$, with $U_{0}$ representing a unit velocity. For simplicity, all simulations had a Prandtl number $\textrm{Pr}=\nu/\kappa=1$. The kinematic viscosity was chosen such that the Kolmogorov scale $\eta=(\nu^{3}/\varepsilon)^{1/4}\approx 0.005L_{0}$ was well resolved, where the kinetic energy dissipation rate is $\varepsilon=\nu\left<|\bm{\omega}|^{2}\right>$ and $\bm{\omega}=\bm{\nabla}\times{\bf u}$ is the vorticity. This results in $\kappa_{\textrm{max}}\eta\approx 1.6$, where $\kappa_{\textrm{max}}=N_{x}/(3L_{0})$ is the maximum resolved wave number when using the $2/3$ rule for dealiasing, ensuring spatially well-resolved simulations [47, 48].
@@
Particles were initialized randomly in a horizontal strip of width $H/5$, centered around $z_{0}=H/2$, and at a time at which the flows had reached a turbulent steady state (i.e., after flow integration for over 60 large-scale turnover times). The initial velocities of the particles matched the fluid velocity at the center of each particle. The particles were one-way coupled, essentially functioning as test particles. They neither collide with each other nor back react into the flow dynamics. In each fluid simulation in table 1 four sets of particles were introduced (see table 2), each containing 25,000 particles, each set characterized by distinct values of $\tau_{p}$. Particles were integrated for more than 15 large-scale turnover times (simulations with particles with $\textrm{St}=6$ were extended for twice this time, as particles with more inertia have a longer relaxation time; see below for a definition). Moreover, for each set of particles, numerical integrations were done solving Eq. (6) with and without the Basset-Boussinesq history term (labed in the following respectively as “w/H” and “w/oH”). This resulted in a cumulative count of 32 particle datasets, each with their corresponding Reynolds, Froude, and Stokes numbers as defined next.
@@
which quantifies the turbulence intensity at the buoyancy scale $L_{b}=2\pi/k_{b}$, with $k_{b}=N/U$, and plays a crucial role in describing the flow behavior. When $\textrm{Rb}\gg 1$ turbulence is strong even in the presence of stratification, while for $\textrm{Rb}\ll 1$ turbulent motions are significantly suppressed by viscosity in each stratified layer. Geophysical flows typically exhibit large values of Rb [49], but computational constraints impose severe restrictions on the range of values of Rb that can be simulated. Previous studies indicate that $\textrm{Rb}>10$ is sufficient for the flow to sustain significant turbulence at small scales [50]. Considering computational limitations and the need to explore parameter space, we consider flows with Rb ranging from 6 to 130. Another relevant length scale to characterize the small scales of stratified turbulence is the Ozmidov scale, $L_{Oz}=2\pi/k_{Oz}$, with $k_{Oz}=(N^{3}/\varepsilon)^{1/2}$. At scales significantly smaller than $L_{Oz}$ we expect the flow to asymptotically recover isotropy. Thus, when Rb is sufficiently large and $L_{Oz}$ is larger than the the Kolmogorov dissipation scale, we can expect small-scale turbulence to be stronger, and to affect the particle dynamics and transport.
@@
where St is the aforementioned Stokes number, $\tau_{\eta}$ is the Kolmogorov dissipation time defined before, and $\textrm{Re}_{p}$ is the particle Reynolds number.
Tables 1 and 2 provide all these dimensionless numbers and characteristic scales for the simulations. Note these tables should be read together, as we can have, e.g., particles with $\textrm{St}=0.3$ in a flow with $N=4/T_{0}$ (with or without the Basset-Boussinesq history term), or the same particles but in a flow with $N=8/T_{0}$, $12/T_{0}$, or $20/T_{0}$. To contextualize the choice of parameters, note that in oceanic flows the Froude number varies from $\textrm{Fr}\approx 10^{-2}$ at large scales, to $\textrm{Fr}\approx 10^{-1}$ at vertical scales of the order of $H=1$ km [51]. Recalling that the typical oceanic Brunt-Väisälä frequency is $N\approx 10^{-3}$ s-1, putting units in simulation N08 with $\textrm{Fr}=0.11$, the values of $N$ and $H=1$ km set $L_{0}=637$ m (i.e., the numerical domain has size $4\times 4\times 2$ km3) and $T_{0}=8000$ s. Using these units, the simulation has r.m.s. horizontal velocities of $\approx 0.08$ m s-1, comparable to typical velocities in the ocean of $0.1$ m s-1 [51]. Stokes numbers considered for the particles in table 2 are also within the typical values discussed for the ocean in Sec. II. Of course, the separation of scales in the simulations is much smaller than in realistic oceanic flows. As a result, for the particles we can only compare their characteristic timescales with those of the smallest dynamical scales of the fluid resolved in the simulation, with the main aim of validating the bounds in Sec. III. Attaining realistic separations of scales for geophysical flows (together with realistic values of the Prandtl number, which is here set to 1) is impossible with current computational capabilities, a limitation of this study that should be kept in mind.
@@ V Numerical results
Figure 1: (a) Ratio of the squared infinity norm of the vertical components of the Basset-Boussinesq force to the Stokes force, as a function of the buoyancy Stokes number Sb for all particles and simulations. A linear relation with slope of 1 is shown as a reference. (b) Same for the infinity norms considering all components of both forces.
@@ V.1 Relation between the forces and the buoyancy Stokes number
The first direct test of the bound given by Eq. (19) considers the strength of the forces in all simulations with particles that include the Basset-Boussinesq force. To estimate $\|{\bf F}_{\textrm{BB}}\|_{\infty}$, $\|{\bf F}_{\textrm{BB}}\cdot\hat{z}\|_{\infty}$, $\|{\bf F}_{\textrm{St}}\|_{\infty}$, and $\|{\bf F}_{\textrm{St}}\cdot\hat{z}\|_{\infty}$, the maximum of the vector modulus and of the vertical component of each force were calculated for each particle in the simulations over all times, and then the average values of ${\|{\bf F}_{\textrm{BB}}\|^{2}_{\infty}}/{\|{\bf F}_{\textrm{St}}\|^{2}_{\infty}}$ and of ${\|{\bf F}_{\textrm{BB}}\cdot\hat{z}\|^{2}_{\infty}}/{\|{\bf F}_{\textrm{St}}\cdot\hat{z}\|^{2}_{\infty}}$ were computed for each value of Sb. The results are shown in Fig. 1. Note that ${\|{\bf F}_{\textrm{BB}}\cdot\hat{z}\|^{2}_{\infty}}/{\|{\bf F}_{\textrm{St}}\cdot\hat{z}\|^{2}_{\infty}}<1$ when $\textrm{St}<1$. The condition on ${\|{\bf F}_{\textrm{BB}}\|^{2}_{\infty}}/{\|{\bf F}_{\textrm{Sb}}\|^{2}_{\infty}}$ being small is also bounded by St, confirming that the vertical component of the Basset-Boussinesq force is the largest, and that it is the component of interest to obtain a bound on the effect of this force on the particles dynamics in stratified flows. Qualitatively similar results are obtained for $L^{2}$ norms of the forces, which are not shown (see however the next subsections for mean squared errors of various quantities with and without the Basset-Boussinesq force).
@@ V.2 Influence of the Basset-Boussinesq force on particle dispersion
where the subindex $i=1,2$ stands for the $x$ or $y$ coordinates, and the supraindex $j$ can be w/H or w/oH (i.e., integration with or without the history term). Figure 2 shows the ratio of the averages over all particles of the mean square horizontal displacements with and without the history term, $\langle{\delta_{x}^{(\textrm{w/H})}}^{2}+{\delta_{y}^{(\textrm{w/H})}}^{2}\rangle/\langle{\delta_{x}^{(\textrm{w/oH})}}^{2}+{\delta_{y}^{(\textrm{w/oH})}}^{2}\rangle$, as a function of time for different Froude and Stokes numbers. Note that in Fig. 2(d), it takes much longer for particles with $\textrm{St}=6$ to reach an extremum and form a small plateau. This is due to the longer relaxation time of these particles, which results in extended transients. As previously mentioned, simulations of particles with $\textrm{St}=6$ in this figure, as well as in the figures below, were extended to twice the duration of the other particles to ensure that the extrema were reached. However, the limitations related to the increasing difficulty in achieving a steady state as St increases must be kept in mind as we move forward.
For small times compared with the Brunt-Väisälä period, $2\pi/N$, the ratio in Fig. 2 remains close to unity, but as time becomes close to $2\pi/N$ differences in the mean squared horizontal dispersion develop. For small values of St (corresponding to small $\tau_{p}$) and large values of Fr (i.e., small values of $N$, but still with $\textrm{Fr}<1$), the ratio hovers around unity and the differences between the particles w/H and w/oH remain below $4\%$. As the value of St of the particles increases, the ratio deviates from unity. For fixed St values, the ratio also increases as Fr decreases. This increase in the differences w/H and w/oH is compatible with the increase in Sb. For large values of Sb differences in the horizontal dispersion are between $10\%$ to $15\%$, depending on the value of Fr. However, when $\textrm{Sb}<1$, errors when neglecting the Basset-Boussinesq force consistently stay below $5\%$. It is noteworthy that in many cases the largest errors are found for intermediate times. Van Aartrijk and Clercx [26] already reported that in stratified flows different regimes develop in the vertical and horizontal dispersion of particles, with the Basset-Boussinesq force having a significant impact in the development and extension of transients at times close to the Brunt-Väisälä period.
@@ V.4 Effect of the Basset-Boussinesq force on particle clustering
The vertical confinement of neutrally buoyant particles in thin layers resulting from stratification has a strong impact on inertial particle clustering [28]. This aggregation of particles is relevant in oceanic flows, in which even the dynamics of large particles are sometimes modeled using empirically modified versions of the Maxey-Riley-Gatignol equation. An example we already discussed can be found in the study of Sargassum, a type of seaweed that serves as the habitat for marine fauna, but can also pose environmental challenges due to its elevated levels of arsenic and heavy metals when decomposing near coastlines [36, 52]. Another example is given by phytoplankton, that aggregates creating intricate structures spanning kilometers [53, 37], and forming thin layers at depths correlated with regions of pronounced fluid density gradients and vertical shear, typically occurring near the bottom of the oceanic mixed layer [54]. We thus quantify now the effect of considering or neglecting the Basset-Boussinesq force on particle clustering. To quantify particle aggregation we use a Voronoï tessellation. This technique has been used before to characterize the preferential concentration of particles in laboratory experiments and in numerical simulations [55, 56, 57, 7, 58, 59, 60, 61].
In a Voronoï tessellation each particle is assigned a cell, corresponding to all the volume (or area, in two-dimensional slices) that is closer to that particle than to any other neighbouring particle. The sizes of the cells are then inversely proportional to the particle density in that region: larger cells correspond to voids (i.e., regions with far apart particles), while smaller cells correspond to clusters (i.e, regions with particles closer to each other). The standard deviation of the volume (or area) of all the Voronoï cells can then we used to quantify the amount of clustering in the flow [55, 56]: when the standard deviation is larger than that of a random Poisson process, there is an excess of clusters and of voids compared with a random homogeneous distribution of particles. While both three- and two-dimensional tessellations have been used to study particle clustering in HIT, in the case of stably stratified flows the confinement of all particles in a layer near a the neutrally buoyant level makes it reasonable to limit the study to two-dimensional tessellations, by projecting all particles into a horizontal plane (see, e.g., [28]).
@@ VI Conclusions
We obtained a bound that is useful to determine under what conditions the Basset-Boussinesq force in the Maxey–Riley-Gatignol equation for inertial particles can be neglected, when the particles are submerged in a stratified flow and are at their neutrally buoyant level. The bound is only valid for sufficiently stratified fluids with $\textrm{Fr}<1$, i.e., it is not valid and should not be applied in the limit $N\to 0$ of isotropic and homogeneous flows. The bound motivated the definition of a buoyancy Stokes number, $\textrm{Sb}=N\tau_{p}$, which is the Stokes number of the particles at the fluid buoyancy scale, i.e., it is the ratio of the particle response time to the Brunt-Väisälä period. For sufficiently small Sb the effect of the Basset-Boussinesq force becomes negligible.
This condition was validated using direct numerical simulations of small particles in stably stratified turbulent flows, exploring parameter space by varying the fluid Froude number and the particles Stokes number. Numerical integration of the particles was also performed considering and neglecting the Basset-Boussinesq force in the Maxey–Riley-Gatignol equation, to quantify differences between these two cases.
Using the numerical simulations we computed the infinity norm of the forces acting on the particles, and showed that the infinity norm of the Basset-Boussinesq force becomes smaller than the infinity norm of the Stokes force when $\textrm{Sb}<1$. Three other observables were considered: single particle dispersion (i.e., the mean squared distance traveled by the particles), the mean squared velocity of the particles, and particle clustering. In all cases, differences between simulations with and without the Basset-Boussinesq force were small when $\textrm{Sb}<1$, and grew rapidly with Sb when $\textrm{Sb}>1$. Thus, the Basset-Boussinesq force must be considered in stratified flows only for particles with very large inertia, or when the stratification is too strong. The results allow the estimation of the conditions under which this force becomes relevant in different geophysical contexts [53, 22, 26, 36, 52, 5, 6]. It is interesting that for typical parameters of oceanic flows, even particles of 50 mm radius have $\textrm{Sb}\approx 0.1$, and thus neglecting the Basset-Boussinesq force may be justified in most cases.